# Quantum Optical Electron Pulse Shaper

### v1

@@ Abstract
Coherent control of ultrafast quantum phenomena benefits from the pulse-shaping capabilities allowing to modulate the envelope and instantaneous phase of optical fields on femtosecond time scales. While such control is available for optical fields, an analogy of a pulse shaper for freely propagating electrons is lacking. In this study, we theoretically demonstrate a method that enables near arbitrary light-based shaping of electron wave packets in the time domain. The method is based on the quantum phase modulation of electron waves by coherent light with time-dependent frequency leading to generation of spectrally separated electron energy side bands with shaped time-energy profiles and envelopes. Our results show that few femtosecond time durations can be achieved without additional spectral broadening of the electron wave packet, allowing one to reach the combination of high time, spatial, and spectral resolutions in ultrafast imaging and diffraction experiments with pulsed electron beams.
@@
I Introduction
Pulsed electron sources have been developed with the aim of extending the exceptional spatial resolution of electron microscopes to the time domain. Ultrafast electron microscopy and diffraction evolved in complex tools enabling observation of dynamical phenomena occurring on nanoscale, such as structural and phase transitions [1, 2], magnetic [3, 4] and charge carrier dynamics [5, 6, 7], atomic motion [8] and plasmonic [9] and phonon [10, 11] dynamics.
Pulsed electron beams for ultrafast electron microscopy are routinely generated via laser-stimulated photoemission. The energy bandwidth of the emitted electron wave packet is determined by the geometry and material properties of the emitter, the applied electric field, and the spectrum of the photoemission light. The nonzero energy spread combined with dispersive propagation of electrons lead to elongation of the generated pulses during their acceleration and propagation to the specimen [12]. As a result, the shortest pulse duration of electrons achieved in electron microscopes without additional compression is on the order of few hundreds of femtoseconds [13, 14, 15]. Such time resolution allows one to study picosecond processes, but the time scales associated, for example, with coherent lattice vibrations or coherent electron dynamics are not directly accessible without additional compression of the electrons.
@@ I Introduction
Although the phase and envelope of ultrashort optical pulses can be manipulated in the time domain by various shaping methods [27, 28], such complex control has not been available for pulsed electron beams. Optical pulse shaping techniques are typically based on separate control of the phase and amplitude of different spectral components of the pulse [28]. Their coherent superposition then determines the time profile of the pulse envelope. However, ultrashort electron pulses differ strongly from their optical counterparts due to the fact that their coherence time is much shorter than the Fourier-limited pulse duration [29]. In contrast to fully coherent optical waves generated by lasers, the electron wave packets from time-separated regions within the same pulse do not coherently add up to shape the envelope of the pulse.
Here we propose a quantum-optical electron pulse shaping concept, which is based on the interaction of a partially coherent electron pulse with shaped optical fields with time-varying frequency. Due to the energy-momentum dependence mismatch between photons and electrons in free space [30], an efficient energy modulation of the electrons requires breaking the spatial inversion symmetry of the optical fields. The light-electron interaction can be facilitated by a nearby plasmonic nanostructure [20, 31, 32, 21, 33, 34, 14], a metallic or dielectric membrane [25, 35, 36, 37, 24], an evanescent optical field [38, 39, 40, 41, 42, 43] or by the ponderomotive interaction with multiple optical beams in vacuum [22, 23, 44]. When the electrons interact with quasi-monochromatic light, they absorb and emit an integer number of photons. The resulting electron energy spectrum consists of sidebands separated by integer multiples of the laser photon energy $\hbar\omega$ [45]. Here we propose to utilize light pulse with time-dependent frequency $\omega(\tau)$. In such case, the distance between the neighboring photon side bands generated in the electron energy spectrum in any instant of time $\tau$ is given by the instantaneous photon energy $\hbar\omega(\tau)$ [46].
By tailoring the optical pulses used for coherent manipulation of the electron wave function, the longitudinal phase space of individual photon sidebands within the electron pulse can be precisely controlled and shaped, constrained only by the bandwidth of the optical pulses and the limits of optical pulse shaping. Combined with subsequent dispersive propagation of electrons to the specimen and with spectrally resolved detection, this technique can be utilized to generate close to arbitrary shape of the electron pulse in the longitudinal phase space. In this work we focus on application of this principle to compression of electron pulses and compensation of nonlinear chirp of the electrons.
II Electron Pulse Shaping Principle
Figure 1: (a) The duration of the electron pulse after the photoemission is determined by the time envelope of the the triggering optical pulse and the spectral width is dependent on the photoemission conditions. (b) After the emission, for distance $z_{\text{el}}$ the electron pulse is accelerated to a desired energy and is let to propagate dispersively to the site of the pulse-shaping interaction. During the process the electron pulse is elongated in time and acquired a positive chirp. At the interaction site the electron pulse traverses a dielectric membrane e.g., illuminated by an optical pulse with modulated instantaneous photon energy $\hbar\omega(\tau)$ and intensity envelope $I(\tau)$. (c) The absorptions and emission of photon quanta result in the generation of tailored energy sidebands, where the instantaneous energy separation between the bands is determined by $\hbar\omega(\tau)$ and the instantaneous quasi-probability $I(\tau)$. (d) After propagating for a sufficient distance, the quesi-probability side bands are reshaped and some of them experience temporal compression, the propagation distance denoted $z_{\text{prop}}$. The final time duration of the selected electron sub-pulse is denoted as $\tau_{\text{com}}$ and in dependence on the light field parameters, can be made shorter than the initial pulse duration by a factor of $M$.
@@ II Electron Pulse Shaping Principle
The stimulated inelastic interaction between a periodic vector potential of light and the electron wave function can be understood in a semi-classical and nonrecoil approximations as a periodic phase modulation of the electron wave function [30]. The final wave function after the interaction can be described as a superposition of discrete momentum states separated by the momenta of the interacting photons. This principle can be generalized to the case of optical fields with time-dependent frequency, provided that the change in light frequency within the electron coherence time is negligible [46].
@@
In the quantum optical electron pulse shaper proposed here, the electron pulse is generated by photoemission from a photocathode, accelerated by a static electric field, and propagates to the plane where pulse-shaping interaction is applied [see the layout in Fig. 1 (a)]. We assume that the duration of an electron pulse immediately after photoemission is determined by the time envelope of the optical pulse incident on the cathode in the electron source [12]. In contrast, the electron coherence time depends on the photoemission conditions. Without additional monochromatization, the typical electron energy spread in an ultrafast transmission electron microscope equipped with a field emission source is $\delta E=$0.3-0.7 eV [13, 47] with the associated coherence time of electrons $\tau_{\text{coh}}=\hbar/\delta E\approx$2-5 fs [12, 29].
@@
In the interaction plane, the electron pulse interacts with a shaped optical pulse at a dielectric membrane [Fig. 1 (b)], ensuring efficient coupling of electrons and photons [48]. The electric field of the optical pulse generated in an optical pulse shaper [27, 28] can be described using an envelope and a time-dependent frequency $\omega(\tau)$. Due to the coherent phase modulation of the electron wave function at the instantaneous frequency of the shaped light pulse, the separation of the generated electron energy sidebands $\hbar\omega(\tau)$ becomes a function of time [Fig. 1 (c)]. Moreover, their instantaneous populations can be controlled by the envelope of the optical pulse. These two virtually independent quantities give us a tool for complex shaping of the electron wave in the longitudinal phase space. Combined with subsequent dispersive propagation between the pulse shaping plane and a sample, the electron pulse can be structured in time domain leading to a single compressed pulse or a periodic train of compressed pulses in one of the electron energy sidebands [Fig. 1 (d)], which can be isolated from the rest of the electron distribution by spectral filtering.
III Theoretical model
$\hat{H}=\frac{1}{2m_{e}}\left(\hat{\mathbf{p}}+e\mathbf{A}\right)^{2}=\hat{H}_{0}+\hat{H}_{\text{int}},$ (1)
where $e>0$ is the elementary charge and $\hat{\mathbf{p}}$ is the electron momentum operator. The free-particle Hamiltonian $\hat{H}_{0}=\hat{\textbf{p}}^{2}/2m_{e}$ describes the evolution of the wavepacket in free space, while the interaction Hamiltonian $\hat{H}_{\text{int}}$ describes the interaction with an electromagnetic field.
III.0.1 Before interaction
$\tilde{\psi}(E)=\mathcal{N}_{E}e^{\left[-\frac{(E-E_{0})^{2}}{2\sigma_{E}^{2}}\right]}e^{\left[-i\frac{\alpha(E-E_{0})^{2}}{2\hbar^{2}}\right]}e^{\left[-i\frac{\beta(E-E_{0})^{3}}{6\hbar^{3}}\right]},$ (2)
where $\mathcal{N}_{E}$ is a normalization constant, $\sigma_{E}=\delta E/\sqrt{4\ln(2)}$ the energy width, $\alpha$ is the group delay dispersion (GDD) coefficient and $\beta$ is the third order group dispersion coefficient (TOD). The time-domain representation is connected to the energy representation via Fourier transform $\psi(\tau)=\mathcal{F}\left\{\tilde{\psi}(E)\right\}$. To account for a partially coherent state, we introduce the density matrix $\rho(\tau,\tau^{*})$
$\rho(\tau,\tau^{*})=\sum_{\tau_{0}}w_{\tau_{0}}\psi(\tau-\tau_{0})\psi^{\dagger}(\tau^{*}-\tau_{0}),$ (3)
$w_{\tau_{0}}=\mathcal{N}_{w}\exp{\left[-\frac{8\ln{2}\tau_{0}^{2}}{\tau_{\text{pe}}^{2}}\right]},$ (4)
In phase space we represent the electron state via its Wigner function $W(E,\tau)$, derived from the density matrix $\rho(\tau,\tau^{*})$ via Wigner-Weyl transformation [49]. To model a dispersion-free state, we set $\alpha=\beta=0$ in Eq. (2), compute the partially incoherent density matrix $\rho_{0}(t,t^{*})$ via Eq. (3) and apply the Wigner transformation to obtain $W_{0}(E,t)$. This state has a probability density envelope FWHM duration of $\tau_{0}=\sqrt{\tau_{\text{coh}}^{2}+\tau_{\text{pe}}^{2}}$. For a dispersed state, we use nonzero $\alpha$ and $\beta$ to get $\rho_{1}(\tau,\tau^{*})$ and $W_{1}(E,\tau)$ with an extended FWHM duration $\tau_{\text{el}}=\sqrt{\tau_{0}^{2}+\tau^{2}_{\text{chirp}}}$, where $\tau_{\text{chirp}}$ accounts for dispersion induced elongation due to propagation between the photocathode and the interaction plane.
III.0.2 Interaction
$\phi(\textbf{r},t)=\left[-\frac{1}{\hbar}\int^{t}_{-\infty}\hat{H}_{\text{int}}(\textbf{r}+\textbf{v}_{0}(\theta-t),\theta)\,\mathrm{d}\theta\right],$ (5)
For a monochromatic electric field with frequency $\omega$, the vector potential is expressed in terms of the electric intensity $\mathbf{A}(\mathbf{r},t)=(-ic/\omega)\mathbf{E}(\mathbf{r},t)+c.c.$. A polychromatic optical incident field with spectrum ~$\tilde{\mathcal{E}}(\omega)=\sqrt{I(\omega)}e^{i\varphi(\omega)}$, where $I(\omega)$ is the spectral intensity and $\varphi(\omega)$ is the spectral phase, can be expressed as a superposition of the monochromatic components
$\textbf{E}_{i}(\mathbf{r},t)=\int_{0}^{\infty}\textrm{d}\omega\,\sqrt{I(\omega)}e^{i\varphi(\omega)}e^{-i\omega(t+\mathbf{s_{i}}\cdot\mathbf{r}/c)},$ (6)
$\textbf{E}_{i}(\mathbf{r},t)\propto\sqrt{I(\mathbf{r},t)}\exp{\left[-i\varphi(\mathbf{r},t)\right]}.$ (7)
Expanding the phase into polynomial series of the wave argument $(t-\mathbf{s}\cdot\mathbf{r}/c)$ we obtain $\varphi(\textbf{r},t)=\omega_{0}(t-\mathbf{s}_{i}\cdot\mathbf{r}/c)+a(t-\mathbf{s}_{i}\cdot\mathbf{r}/c)^{2}+...$, where $\omega_{0}$ is the central angular frequency and $a$ is the linear chirp coefficient.
$\textbf{E}(\mathbf{r},t)=\int_{0}^{\infty}\textrm{d}\omega\,\sqrt{I(\omega)}e^{i\varphi(\omega)}\mathbf{f}_{0}(\mathbf{r},\omega)e^{-i\omega t},$ (8)
where $\mathbf{f}_{0}(\mathbf{r},\omega)$ contains the spatial dependence of the field on frequency $\omega$ [48, 35].
For further evaluation, we consider 1D propagation of the electrons along $z$ with momentum vector $\mathbf{p}_{0}=(0,0,p_{0})$, group velocity vector $\mathbf{v}_{0}=(0,0,v_{0})$ and we only take into account the $z$ component of the electric field $f_{0z}$. Plugging into Eq. (8) into Eq. (5), we evaluate the evolution operator:
$\begin{split}U(z,t)=\\ \exp{\left\{-i2\text{Im}\int_{0}^{\infty}\mathrm{d}\omega\,g(\omega)\sqrt{I(\omega)}e^{i\varphi(\omega)}e^{-i\omega(t-z/v_{0})}\right\}},\end{split}$ (9)
$g(x,y,\omega)=\frac{e}{\hbar\omega}\int_{-\infty}^{z}\mathrm{d}z^{\prime}\,f_{0z}(x,y,z^{\prime};\omega)e^{i\omega z^{\prime}/v_{0}}.$ (10)
$U(\tau)=\exp{\left\{-i2|g|\sqrt{I(\tau)}\sin{[\varphi(\tau)+\arg{g}]}\right\}}.$ (11)
In practice, we shift all the factors to $g$, so that $I(\tau)$ is normalized to 1 at the maximum. The density matrix describing the state of the electrons after the interaction is:
$\rho_{2}(\tau,\tau^{*})=U(\tau)\rho_{1}(\tau,\tau^{*})U^{\dagger}(\tau^{*}),$ (12)
III.0.3 After interaction
As a result of the interaction, sidebands with shaped time-dependent energy evolution are generated in the Wigner function $W_{2}(E,\tau)$, which are expected to be altered by dispersive propagation in vacuum for time $\theta_{\text{prop}}$. We emphasize that $\tau$ is the detection time within the electron pulse, whereas $\theta$ is the running evolution time. The evolution is described as coordinate transformation governed by the free-electron Hamiltonian $\hat{H}_{0}$ with the Hamilton equation of motion derived from its classical analogue $H_{0}$. The new coordinates after propagation time $\theta_{\text{prop}}$ are
$\tau^{\prime}-\tau^{\prime}_{0}=\tau-\tau_{0}+\frac{[v(E)-v_{0}]\theta_{\text{prop}}}{v_{0}},$ (13)
$E^{\prime}-E_{0}^{\prime}=E-E_{0},$ (14)
where $v[E(p)]=\partial E(p)/\partial p$ is the momentum-dependent group velocity. The transformed Wigner function is $W_{3}(E,\tau,\theta_{\text{prop}})=W_{2}(E^{\prime},\tau^{\prime})$.
IV Results
@@
The pulse shaping method is examined in two model situations for lower kinetic energy electrons (5 keV) and higher kinetic energy electrons (30 keV). In the following section, we show the electron pulse represented by a non-negative spectrogram $S(E,\tau)$ defined as a convolution of the Wigner function with a Gaussian kernel
@@
where $\tau_{s}\approx 5\,\mathrm{fs}$. This representation allows us to better visualize the energy band populations and smooths out the fast oscillating coherences between the bands [46].
IV.1 Chirp Inversion Induced Pulse Compression
@@
We illustrate the temporal compression principle on the example of low energetic electrons with central kinetic energy of $5\,\mathrm{keV}$. Low-energy electrons, in general, undergo significant temporal elongation due to dispersion already after short distances, reaching durations several times longer than the initial length of the pulse. Also, the compression of the low energetic electrons can be conveniently achieved in relatively short propagation lengths.
@@
We model the chirp-free electron pulse ($\alpha=0=\beta$ in Eq. (2)), with coherence time $\tau_{\text{coh}}=3.65\,\mathrm{fs}$, energy spread FWHM of $\delta E=0.5\,\mathrm{eV}$ and photoemission pulse duration $\tau_{\text{pe}}\approx 9.30\,\mathrm{fs}$, yielding an electron pulse with total FWHM duration of $\tau_{0}=9.9\,\mathrm{fs}$ [Fig. 2 (a)]. The central group velocity is $v_{0}=0.14c$.
We introduce second order group delay dispersion term only $\alpha=328\,\mathrm{fs^{2}}$, resulting in elongation of the pulse to $\tau_{\text{el}}=250\,\mathrm{fs}$ FWHM. The spectrogram of the chirped electron pulse is shown in [Fig. 2 (b)].
@@
The spectral domain representation of the optical pulse we use for chirp inversion is explicitly
@@
~$\tilde{\mathcal{E}}(\omega)\propto\exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}\right],$ (16)
@@
where $\omega_{0}=1.8\,\mathrm{fs^{-1}}$ is the central photon frequency, the spectral width is matched to that of the electron wave packet $\sigma_{\omega}=\sigma_{E}/\hbar$ and the opposite GDD is used $\varphi^{\prime\prime}=-\alpha$. We consider constant coupling strength of $g=1.75$ to maximize the population of the chirp-inverted sideband.
The spectrogram [Fig. 2 (c)] of the electron pulse after the interaction consists of the ladder of energy bands spaced by integer multiples of the photon energies. This choice of instantaneous photon energy leads to the generation of a sideband with an inverted chirp with respect to the original electron wavepacket $E(\tau)+2\hbar\omega(\tau)$,centered around $E_{0}+2.4$ eV, corresponding to absorption of two photons. Around $20\%$ of the electron probability distribution is transferred to this band.
@@
Under these conditions ($E_{0}\gg\hbar\omega$, $[v(E)-v_{0}]\propto E$), we can derive the required propagation time $\theta_{\text{com}}$, after which the side-band will compress. If the central group velocity of the second sideband is $v_{2\hbar\omega_{0}}=v(E_{0}+2\hbar\omega_{0})$, then the electron pulse needs to propagate for $z_{\text{com}}=\frac{m_{e}v_{2\hbar\omega}^{2}\alpha}{\hbar}\approx 21\,\mathrm{cm}$. We note this position as the compression point [Fig. 2 (d), (e)]. The side-band is compressed down to the initial duration of the pulse (numerically evaluated $\approx 9.9\,\mathrm{fs}$) reversing the elongation while maintaining the initial spectral width.
@@
Figure 2: Spectrogram representation of the electron pulse (a) dispersion-free, with total pulse duration of 9.9 FWHM, (b) with group delay dispersion $\alpha$ and total duration of 250 fs FWHM. Spectrogram representation (c) of the electron side-bands after interaction with the optical field, (d) after propagation, (e) close-up of the compressed sideband spectrogram. We note that in the spectrogram representation the sidebands are elongated in the $\tau$-direction due to the convolution with Gaussian kernel.
IV.2 Nonlinear Electron Chirp Correction
Next we demonstrate the correction of higher order dispersion by introducing nonzero TOD to the electron wave. The parameters $E_{0}$, $\Delta E$ and $\alpha$ remain the same as in the previous case. The TOD is $\beta=256\,\mathrm{fs^{3}}$. The overall FWHM duration of the electron pulse is $\tau_{\text{el}}=242$ fs.
@@ IV.2 Nonlinear Electron Chirp Correction
~$\begin{split}\tilde{\mathcal{E}}(\omega)\propto\\ \exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}-i\frac{\varphi^{\prime\prime\prime}(\omega-\omega_{0})^{3}}{6}\right]\end{split}$ (17)
The curvature is compensated in the second side-band by a light pulse with the same GDD as previously and TOD $\varphi^{\prime\prime\prime}=-\beta/2=128\,\mathrm{fs^{3}}$.
@@
In Fig. 3 we show the spectrogram of an electron pulse with a combination of linear and nonlinear chirp. In [Fig. 3(a)] interaction with light field with $\varphi^{\prime\prime}=-\alpha$ and $\varphi^{\prime\prime\prime}=0$ is applied. In [Fig. 3(b)] interaction with light field with $\varphi^{\prime\prime}=-\alpha$ and $\varphi^{\prime\prime}=-\beta/2$ is applied. The difference between the two is subtle immediately after the interaction. After further propagation for 21 cm it is noticeable, that the sideband generated with linearly chirped light cannot compress ideally (c) due to the leftover curvature of the band, whereas introducing the $\varphi^{\prime\prime\prime}$ term leads to better compression (d).
IV.3 Short Pulse Gating

### v2

@@ Abstract
Coherent control of ultrafast quantum phenomena benefits from pulse-shaping capabilities allowing to modulate the envelope and instantaneous phase of optical fields on femtosecond time scales. While such control is available for optical fields, an analogy of a pulse shaper for freely propagating electrons is lacking. In this study, we theoretically demonstrate a method that enables near arbitrary light-based shaping of electron wave packets in the time domain. The method is based on the quantum phase modulation of electron waves by coherent light with time-dependent frequency leading to generation of spectrally separated electron energy side bands with shaped time-energy profiles and envelopes. Our results show that few femtosecond time durations can be achieved without additional spectral broadening of the electron wave packet, allowing one to reach the combination of high time, spatial, and spectral resolutions in ultrafast imaging and diffraction experiments with pulsed electron beams.
@@
Introduction—Pulsed electron sources have been developed with the aim of extending the exceptional spatial resolution of electron microscopes to the time domain. Ultrafast electron microscopy and diffraction evolved in complex tools enabling observation of dynamical phenomena occurring on nanoscale, such as structural and phase transitions [22, 7], magnetic [3, 41] and charge carrier dynamics [53, 36, 43], atomic motion [30] and plasmonic [31] and phonon [16, 54] dynamics.
Pulsed electron beams for ultrafast electron microscopy are routinely generated via laser-stimulated photoemission. The energy bandwidth of the emitted electron wave packet is determined by the geometry and material properties of the emitter, the applied electric field, and the spectrum of photoemission light. The nonzero energy spread combined with the dispersive propagation of electrons lead to elongation of the generated pulses during their acceleration and propagation to the specimen [2]. As a result, the shortest pulse duration of electrons achieved in electron microscopes without additional compression is on the order of few hundreds of femtoseconds [8, 26, 34]. Such time resolution allows one to study picosecond processes, but the time scales associated, for example, with coherent lattice vibrations or coherent electron dynamics are not directly accessible without additional compression of the electrons.
@@
Although the phase and envelope of ultrashort optical pulses can be manipulated in the time domain by various shaping methods [50, 51], such complex control has not been available for pulsed electron beams. Optical pulse shaping techniques are typically based on separate control of the phase and amplitude of different spectral components of the pulse [51]. Their coherent superposition then determines the time profile of the pulse envelope. However, ultrashort electron pulses differ strongly from their optical counterparts due to the fact that their coherence time is much shorter than the pulse duration [46]. In contrast to fully coherent optical waves generated by lasers, the electron wave packets from time-separated regions within the same pulse do not coherently add up to shape the envelope of the pulse. While application of phase-modulated laser pulses to control spectral content of X-ray photons emitted by Thomson scattering has been proposed [14, 45], complex temporal shaping of the quantum state of the electrons by temporally modulated optical pulses has not been considered yet.
Electron Pulse Shaping Principle—In this study, we propose a quantum-optical electron pulse shaping concept based on the interaction of a partially coherent electron pulse with shaped optical fields of time-varying frequency.
@@
The stimulated inelastic interaction between a periodic vector potential of quasi-monochromatic light and the electron wave function can be understood in a semi-classical and nonrecoil approximations as a periodic phase modulation of the electron wave function [12]. Efficient modulation of electron energy can be enabled using plasmonic nanostructures [9, 38, 11, 39, 17, 44, 26], metallic or dielectric membranes [48, 49, 23, 10, 33], evanescent optical fields [4, 35, 24, 6, 18, 21], or ponderomotive coupling with multiple optical beams in vacuum [27, 25, 47]. The final wave function after the interaction can be described as a superposition of discrete energy sidebands separated by the energies of the interacting photons $\hbar\omega$ [40], where $\omega$ typically represents a time-independent frequency of the optical field.
Our proposal is to generalize this principle to the case of optical fields with time-dependent photon energy $\hbar\omega(\tau)$, allowing to directly control the immediate sideband separation in phase space. Here we assume that the optical frequency $\omega(\tau)$ changes slowly compared to the electron coherence time. See Supplemental Materials (SM) [1] for detailed derivation of the effect of the light field with time-dependent frequency on the electrons.
@@
The electron pulse is generated by photoemission from a photocathode [Fig. 1 (a)]. We assume that the duration of an electron pulse immediately after photoemission is determined by the time envelope of the optical pulse incident on the cathode in the electron source [2]. In contrast, the electron coherence time depends on the photoemission conditions. Without additional monochromatization, the typical electron energy spread in an ultrafast transmission electron microscope equipped with a field emission source is $\delta E=$0.3-0.7 eV [8, Schröder_2024] with the associated coherence time of electrons $\tau_{\text{coh}}=\hbar/\delta E\approx$2-5 fs [2, 46]. The electron pulse is then accelerated by a static electric field, and afterwards propagates freely to the pulse-shaping plane, while both of the processes contribute to temporal elongation and chirping of the electron wavepacket. We consider an electron wavepacket with less than one electron per pulse, thus we neglect possible space charge effect on the electron model properties and propagation.
@@
In the pulse shaping plane, the electron pulse interacts with a shaped laser pulse at a dielectric membrane [Fig. 1 (b)], ensuring efficient coupling of electrons and photons [32]. The electric field of the laser pulse generated in an optical pulse shaper [50, 51] can be described using an envelope and a time-dependent frequency $\omega(\tau)$. Due to the coherent phase modulation of the electron wave function at the instantaneous frequency of the shaped light pulse, the separation of the generated electron energy sidebands $\hbar\omega(\tau)$ becomes a function of time $\tau$ [Fig. 1 (c)]. Moreover, their instantaneous populations can be controlled by the envelope of the optical pulse. These two virtually independent quantities give us a tool for complex and general shaping of the electron wave in the longitudinal phase space.
Combined with subsequent dispersive propagation between the pulse shaping plane and a sample, the electron pulse can be structured in time domain leading to a single compressed pulse or a periodic train of compressed pulses in one of the electron energy sidebands [Fig. 1 (d)]. Thanks to the sufficient separation between the sidebands, the shaped sideband can be isolated from the rest of the electron distribution by spectral filtering.
Figure 1: (a) An electron pulse with finite spectral width is generated via short pulse emission from the tip. (b) The electron pulse acquires a positive chirp through dispersive propagation. At the interaction site the electron pulse traverses a dielectric membrane e.g., illuminated by an optical pulse with modulated instantaneous photon energy $\hbar\omega(\tau)$ and intensity envelope $I(\tau)$. (c) The absorptions (emission) of photon quanta result in the generation of tailored energy sidebands. (d) After propagating for a certain distance, the quesi-probability side bands are reshaped.
Results—The pulse shaping method is examined in two model situations for lower kinetic energy electrons (5 keV) and higher kinetic energy electrons (30 keV). In each special case we aim to reduce the duration of the pulse or pulses in the specimen plane to few fs.
@@
In the following section, we show the electron pulse in energy-time phase space represented by a non-negative spectrogram $S(E,\tau)$ defined as a convolution of the Wigner function $W(E,\tau)$ of the electron wave-packet with a Gaussian kernel
@@
where $\tau_{s}\approx 5\,\mathrm{fs}$. By definition the Wigner function is allowed to exhibit negative values. We opt for spectrogram representation, because it allows us to better visualize the energy band populations and smooths out the fast oscillating coherences, which occur between the bands [42]. Detailed description of the theory and numerical simulations can be found in SM [1].
@@
Chirp Inversion Induced Pulse Compression— A special case of pulse shaping is temporal compression of electron pulses induced by chirp inversion in a single energy sideband. We use as an example low energetic electrons with central kinetic energy of $5\,\mathrm{keV}$. Low-energy electrons, in general, undergo significant temporal elongation due to dispersion already after short distances, reaching durations several times longer than the initial length of the pulse. Also, the compression of the low energetic electrons can be conveniently achieved at relatively short propagation distances.
@@
After photoemission, the electron wavepacket has an envelope with duration $\tau_{0}=(\tau_{\text{coh}}^{2}+\tau_{\text{pe}}^{2})^{1/2}=9.9\,\mathrm{fs}$, with coherence time $\tau_{\text{coh}}=3.65\,\mathrm{fs}$, photoemission pulse duration $\tau_{\text{pe}}\approx 9.30\,\mathrm{fs}$ and spectral width $\delta E=0.5\,\mathrm{eV}$ [Fig. 2 (a)]. We consider acceleration to $5\,\mathrm{keV}$ (central group velocity $v_{0}=0.14c$) and dispersive propagation, resulting in elongation to total duration of $\tau_{\text{el}}=250\,\mathrm{fs}$, due to second order group delay dispersion (GDD) term $\alpha=328\,\mathrm{fs^{2}}$ [Fig. 2 (b)].
@@
The spectral domain representation of the optical pulse used for chirp inversion is explicitly
@@
$\tilde{\mathcal{E}}(\omega)\propto\exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}\right],$ (2)
@@
where $\omega_{0}=1.8\,\mathrm{fs^{-1}}$ is the central photon frequency, the spectral width is matched to that of the electron wave packet $\sigma_{\omega}=\delta E/\hbar\sqrt{4\ln(2)}$, the energy width, and the opposite GDD coefficient is used $\varphi^{\prime\prime}=-\alpha$.
The spectrogram [Fig. 2 (c)] of the electron pulse after the interaction with light consists of the ladder of energy bands spaced by integer multiples of the instantaneous photon energy. We consider constant coupling strength of the interaction $g=1.75$ (see [1] for definition), so about 20% of the total electron probability distribution is transferred to the second-order sideband, where the reversal of the initial chirp is achieved.
@@
We can derive the required propagation time $\theta_{\text{com}}$ for the second side-band in the electron probability density to compress. We consider that the photon energy is small compared to the central electron energy $E_{0}\gg\hbar\omega$ and the group velocity $v(E)$ in the vicinity of $v_{0}$ varies linearly with the energy $v(E)-v_{0}]\propto E$. With central group velocity of the second sideband $v_{2\hbar\omega_{0}}=v(E_{0}+2\hbar\omega_{0})$, the electron pulse needs a propagation distance of $z_{\text{com}}=\frac{m_{e}v_{2\hbar\omega}^{2}\alpha}{\hbar}\approx 21\,\mathrm{cm}$. We label this position in $z$ as the compression point. The second sideband reaches the initial duration of the pulse (numerically evaluated $\approx 9.9\,\mathrm{fs}$) reversing the elongation while maintaining the initial spectral width. [Fig. 2 (d), (e)].
@@
Figure 2: Spectrogram representation of the electron pulse (a) dispersion-free, with total pulse duration of 9.9 FWHM, (b) with group delay dispersion $\alpha$ and total duration of 250 fs FWHM, (c) the electron side-bands after interaction with the optical field, (d) after propagation, (e) close-up of the compressed sideband spectrogram. We note that in the spectrogram representation the sidebands are elongated in the $\tau$-direction due to the convolution with Gaussian kernel.
Nonlinear Electron Chirp Correction—To demonstrate the capabilities of electron pulse shaping for correction of higher order dispersion, we add a third order dispersion term (TOD) $\beta=256\,\mathrm{fs^{3}}$ to the electron pulse with the same parameters as used in the calculations presented in Fig. 2. In this case the electron pulse phase space representation is curved and more importantly, correction with linearly chirped light is not sufficient.
In Fig. 3 we show the spectrogram of an electron pulse with a combination of linear and nonlinear chirp. In Fig. 3(a), optimally linearly chirped light is used for the correction. After propagation for 21 cm it is noticeable, that the sideband generated with linearly chirped light cannot compress ideally, reaching 13 fs FWHM duration [Fig. 3(c)] due to the leftover curvature of the band.
@@
$\tilde{\mathcal{E}}(\omega)\propto e^{\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}_{\omega}}-i\frac{\varphi^{\prime\prime}(\omega-\omega_{0})^{2}}{2}-i\frac{\varphi^{\prime\prime\prime}(\omega-\omega_{0})^{3}}{6}\right]}$ (3)
The curvature of the zero-loss band is compensated in the second side-band by a light pulse with the same GDD as previously and TOD $\varphi^{\prime\prime\prime}=-\beta/2=128\,\mathrm{fs^{3}}$, as shown in Fig. 3(b). Further propagation leads to better compression to 11 fs FWHM Fig. 3(d).